The driving distance for professional golfer, Lion Woods, varies uniformly between 250 and 275 yards.

(Round probabilities to four decimals and yardage to two decimals)

a) What's the probability that one of his randomly selected drives is between 252 and 255 yards?

b) What's the 36th percentile of the distribution of his driving distances?

c) What's the standard deviation of his driving variance?

d) 48% of his drives are longer than a specific distance. Find that specific distance.

e) Approximately 68% of his driving distances fall between
Answer #1

ぶ司ニ、/25 2.50 252 250 245 275 250 (a 25S 2 S 251 弒 f(X) d. = 0 , 36 36 2 25d 0-36 25o 259 275 r-2 2Now b 295 , a-250 Vartx) -275 -250 25 | 2、083 =11, 44 31 Aase nce Spect ttc 275 275 2 5 275- K275- 263 2 63 d 68 0 25D m 27SLat 5 div listance. em m and, 25D 2 5 m=250 + 4-254 Aaain, 0 16 40- 6& = o.gq 250 Pso n-5o84 or , n-2 50 89 x25-21 の2 2구! 6